# Beyond Prompts: Space–Time Decoupling Control-Plane Jailbreaks in LLM Structured Output

### v2

@@ 5.2. CDAs Can Effectively Bypass Prompt Guardrails
    EnumAttack on prompt-guarded models. We test open-weight and proprietary LLMs under a state-of-the-art prompt guard, JBShield (JBShield2025usenixsecurity), using AdvBench as the task suite; we report ASR and guard detection rates.
@@
    Comparison to prompt-based jailbreaks. Using llama-3.1-8b, we reproduce five representative prompt attacks (jailbreak2024ccs; drattack2024emnlpfinding; puzzler2024aclfinding; zulu2024arxiv; base64attack2023nips) under the same detection setting as JBShield (JBShield2025usenixsecurity), contrasting control-plane (schema) vs. data-plane (prompt) vulnerabilities.
@@
    Infrastructure sensitivity. We ablate serving stack and grammar backend, covering two serving frameworks (vLLM and SGLang) and two grammar engines ([xgrammar] and [outlines]). We measure whether CDA effectiveness is tied to a particular implementation.
@@
Table 2. EnumAttack ASR on open-weight and proprietary LLMs on AdvBench. Results w/ and w/o prompt guardrail are listed outside/inside parentheses accordingly.
Model ASR
w/ (w/o)
phi-3.5-moe (phi3safety2024arxiv) 99.2% (99.2%)
mistral nemo (mistralnemo2024web) 100.0% (100.0%)
qwen-2.5-32b (qwen2.5_2024arxiv) 99.0% (99.0%)
llama-3.1-8b (llama3-2024website) 99.2% (99.2%)
gemma-2-9b (gemma2_2024arxiv) 98.6% (98.6%)
deepseek-v3 (deepseekv3-2025arxiv) 98.7% (98.7%)
gpt-4o (gpt4o2024web) 100.0% (100.0%)
gpt-4o-mini (gpt4omini2024web) 99.0% (99.0%)
gemini-2.0-flash (gemini2023arxiv) 95.8% (95.8%)
As shown in Table 2, EnumAttack achieves near-100% ASRs across all evaluated open-weight and proprietary models protected by JBShield. This outcome underscores EnumAttack’s high effectiveness to break existing prompt-based guardrails.
5.2.2. Comparison to prompt-based jailbreaks.
Table 3. ASR(%) on llama-3.1-8b with JBShield (JBShield2025usenixsecurity). While other attacks (jailbreak2024ccs; drattack2024emnlpfinding; puzzler2024aclfinding; zulu2024arxiv; base64attack2023nips) are effectively mitigated, JBShield is fully bypassed by our EnumAttack, which achieves even higher ASR by embedding the malicious query in the JSON schema payload.
Method IJP(jailbreak2024ccs) DrAttack(drattack2024emnlpfinding) Puzzler(puzzler2024aclfinding) Zulu(zulu2024arxiv) Base64(base64attack2023nips)
@@
As depicted in Table 3, JBShield detects prompt-based attacks with very high accuracy, leading to near 0% ASR regardless of the original attack’s performance, demonstrating effectiveness against traditional data-plane threats. However, JBShield fails to detect EnumAttack, achieving identical results with/without guard. This discrepancy arises because JBShield is a lightweight defense that relies on static attack profiles and model-specific white-box states of the prompt. By maintaining a completely benign prompt body, CDAs like EnumAttack completely bypass its detection mechanism, rendering such static, data-plane-only defenses inapplicable to control-plane vulnerabilities. Notably, the lengthy question body in IJP dataset (jailbreak2024ccs) doesn’t fit well in the template of EnumAttack, leading to 47.8% ASR in one query. While the judge model ft-mistral-sorrybench-2024-06 (sorrybench2025iclr) doesn’t know Zulu and Base64 languages, and the ASRs are therefore not accurately assessed, leading to lower ASR, which is an inherent limitation for the LLM-as-a-judge method during jailbreak evaluation. As a result, we use gpt-4o as a more powerful and accurate LLM judge for all other evaluations as stated in §§ 5.1.
@@
Figure 5. EnumAttack evaluation of using different grammar backends and serving engines with llama-3.1-8b, except negligible differences caused by temperature and minor variations in the backend, the attack is consistently successful.
To ensure our findings are not backend-specific, we evaluate two grammar engines (xgrammar (xgrammar2024arxiv) and outlines (outlines2023arxiv)) across mainstream serving stacks (vllm (vllm2023sosp) and sglang (sglang2024arxiv)). As shown in Figure 5, EnumAttack remains consistently successful across all combinations, demonstrating that the vulnerability is inherent to the mechanism rather than a specific implementation flaw.
@@
Table 4. Major ASR results on mainstream deployed models (gpt-4o, gpt-4o-mini, gemini-2.0-flash). Results w/ and w/o grammar guardrail are listed outside/inside parentheses accordingly. We use llama-guard-3-8b as the grammar guardrail and set k=1k{=}1 for DictAttack.
gpt-4o Baseline 1.2%(1.2%) 26.0%(26.0%) 10.0%(10.0%) 33.9%(33.9%) 5.1%(5.1%)
EnumAttack 3.4%(100.0%) 7.0%(100.0%) 8.0%(100.0%) 16.6%(96.4%) 5.1%(99.4%)
DictAttack 98.1%(99.8%) 98.3%(100.0%) 98.3%(100.0%) 96.9%(98.6%) 96.7%(98.4%)
gpt-4o-mini Baseline 2.1%(2.1%) 44.0%(44.0%) 14.0%(14.0%) 42.5%(42.5%) 8.9%(8.9%)
EnumAttack 3.4%(99.0%) 7.0%(100.0%) 8.0%(97.0%) 16.8%(96.4%) 5.1%(98.1%)
gemini-2.0-flash Baseline 17.5%(17.5%) 46.0%(46.0%) 22.0%(22.0%) 29.3%(29.3%) 6.0%(6.0%)
EnumAttack 3.3%(95.8%) 7.0%(92.0%) 8.0%(92.0%) 15.5%(89.5%) 4.2%(81.2%)
DictAttack 96.8%(98.5%) 97.3%(99.0%) 89.5%(91.0%) 93.9%(95.5%) 91.4%(93.0%)
Model Adv Harm JB Sorry SR
5.3. Evaluation against Grammar-guarded Systems
In this section, we systematically evaluate the effectiveness of CDAs against grammar-based guardrails across two dimensions:
  1. (i)
  2. (ii)
5.3.1. Comparative Evasiveness: CDA vs. SOTA Data-Plane Attacks
To further validate the robustness of our selected guardrails and highlight the unique threat posed by CDAs, we benchmark the fundamental evasiveness of DictAttack against well-known SOTA data-plane jailbreaks: PAIR (PAIR2024arxiv), TAP (TAP2024nips), and AutoDAN-Turbo (autodanturbo2025iclr).
We conducted this evaluation on gpt-4o using the AdvBench dataset under two settings:
  • Undefended: Direct attack against the model’s internal alignment.
  • Guarded: Data-plane attacks use the best baseline guard; DictAttack is evaluated with our Dual-Plane Guard enabled.
@@ 5.3. Evaluation against Grammar-guarded Systems
The results in Table 6 reveal a critical disparity. Powerful traditional attacks like AutoDAN-Turbo, while highly effective against undefended models (96.0In sharp contrast, DictAttack not only achieves near-perfect success undefended (99.8This underscores that while our system guardrails are effective against known data-plane threats, they are fundamentally insufficient against the cross-plane decoupling of DictAttack. A more detailed analysis of DictAttack’s performance across different configurations is provided in §§§ 5.3.2.
@@ 5.3.2. EnumAttack and DictAttack on grammar-guarded models
Table 4 and Table 5 report ASRs across five benchmarks. The baseline uses each benchmark’s original prompts, while EnumAttack and DictAttack wrap the same intents using our schema-based approach. For each entry, the value outside parentheses is the ASR with the grammar guard enabled and the value in parentheses is the ASR without the grammar guard.
First, baseline results confirm that direct prompting is largely ineffective against the safety alignment of modern LLMs. Although different benchmarks cover different aspects of harmful content, the overall baseline ASR remains consistently low, with averages ranging from 6.7% to 38.7% across datasets. These measurements validate the effectiveness of conventional safety alignment against plain jailbreak attacks.
Second, while EnumAttack achieves an average ASR of 95.8% undefended, simple grammar auditing reduces its success rate to 7.9% (Table 4). This confirms that its reliance on explicit literals makes it highly detectable, which motivates the need for more evasive CDAs like DictAttack.
Third, DictAttack remains highly effective even under grammar auditing. The same grammar-level defense is almost entirely ineffective against DictAttack. As shown in Table 4 and Table 5, DictAttack maintains high ASRs across both mainstream deployed models and flagship/high-capability models. For example, averaged across the five benchmarks, DictAttack achieves 94.3–99.5% ASR on gpt-5, gemini-2.5-pro, deepseek-r1, and gpt-oss-120b without grammar guard (inside parentheses); applying llama-guard-3-8b with k=1k{=}1 (1.7% detection rate) yields 92.7–97.8% ASR with guard (outside parentheses). This demonstrates that the vulnerability is not a flaw of older checkpoints but a fundamental issue that persists in the latest, most heavily aligned models. The slightly lower ASR for gpt-4o-mini (86.4%) is attributed to its reduced reasoning capability, which sometimes fails to correctly reconstruct the query from the dictionary rather than any safety-related refusal.
@@ 5.3.3. DictAttack with combined prompt- and grammar-guarding
Thus, we further evaluate a Dual-Plane Guard setting, adapted from the SOTA SelfDefend (selfdefend2025security). We adapt SelfDefend’s audit prompt to take both the data-plane prompt (keys) and the control-plane grammar (dictionary), using gpt-4o as the auditing LLM.
@@
First, the results demonstrate that even a unified audit fails to effectively mitigate DictAttack. While llamaguard can detect 28.6% of attacks, Moderation APIs almost detect none, probably due to restricted model capabilities. More strikingly, the state-of-the-art SelfDefend (selfdefend2025security)—even with its best strategy—only achieves an 18.5% detection rate against DictAttack on AdvBench. This failure occurs because, while an LLM judge is theoretically capable of reasoning, modern structured output schemas are highly complex and difficult for even powerful LLMs to “parse” and audit semantically in a unified way. The spatial separation of the payload across planes creates a reasoning gap that current auditing-based defenses cannot reliably bridge.
Second, increasing the synthesis factor $k$ markedly weakens the detection. Under Defense 1, gpt-4o attains $\approx 18\%$ detection at $k=0$; however, when $k$ is raised to 10, detection drops to 5.3%, as the malicious tokens are further diluted within the benign dictionary. This reveals that DictAttack exploits a fundamental blind spot: the lack of cross-plane coordination in LLM safety architectures.
Table 7. Detection rate on AdvBench (advbench2023emnlp) under Defense 1 (grammar-only) and Defense 2 (Grammar + Prompt Dual Guard, adapted from SelfDefend; gpt-4o auditor).
@@ 5.3.4. Discussion: Space-Time Decoupling via Interleaved DictAttack
In this variant, the attacker splits the two DictAttack payloads across turns: the prompt payload (keys) is sent first and stored in the conversation context (KV cache), while the grammar payload (dictionary) arrives later in a separate request. Because practical dual-plane guards are typically implemented as per-request auditors, this temporal decoupling effectively reduces a dual-plane audit to a single-plane check, yielding only 5.3% detection for gpt-4o. Mitigating this would require cross-request retrieval and auditing of the entire conversation history at every turn, which is impractical:
  • Unacceptable Cost: Auditing the full history at every turn would require tremendous computational overhead, as the guard system doesn’t know a priori when a malicious payload will arrive.
  • Reasoning Complexity: As the conversation grows (e.g., up to gpt-4o’s 128k token limit), the guard model would require context length and reasoning capabilities comparable to or greater than the protected model itself, which is prohibitively expensive for a real-time monitor.
  • Interpretability: Modern JSON schemas are often dense and opaque. Expecting a guard LLM to semantically reconstruct a query from a dictionary buried 50 turns back in history represents a “needle in a haystack” problem that current systems are not designed to solve.
In conclusion, DictAttack demonstrates that by decoupling intent across both space (planes) and time (requests), adversaries can render external auditing defenses heavy, expensive, and ultimately impractical.

### v3

@@ 5.2. CDAs Open an Attack Surface Orthogonal to Data-Plane Jailbreaks
    EnumAttack on prompt-guarded models. We test open-weight and proprietary LLMs under a state-of-the-art prompt guard, JBShield (Zhang et al., 2025), using AdvBench as the task suite; we report ASR and guard detection rates.
@@
    Compare to prompt-based jailbreaks. On llama-3.1-8b, we reproduce five representative prompt attacks (IJP (Shen et al., 2024), DrAttack (Li et al., 2024a), Puzzler (Chang et al., 2024), Zulu (Yong et al., 2024), Base64 (Wei et al., 2023a)) under the same JBShield (Zhang et al., 2025) guard, contrasting control-plane (schema) vs. data-plane (prompt) vulnerabilities.
@@
    Infrastructure sensitivity. We ablate serving stack and grammar backend, covering two serving frameworks (vLLM and SGLang) and two grammar engines (xgrammar and outlines), to confirm CDA effectiveness is not tied to a particular implementation.
@@
EnumAttack achieves 95.8–100% ASR on AdvBench across all nine models we tested under the JBShield prompt guardrail — five open-weight (phi-3.5-moe, mistral nemo, qwen-2.5-32b, llama-3.1-8b, gemma-2-9b), one large open model (deepseek-v3), and three proprietary models (gpt-4o, gpt-4o-mini, gemini-2.0-flash). The guardrail does not change ASR in any case, since the prompt body remains benign while the malicious intent is hidden inside the JSON Schema; Table 2 below makes this contrast explicit by reproducing five prompt-based attacks under the same guard.
5.2.2. Compare to prompt-based jailbreaks.
To make this asymmetry explicit, we run five representative prompt-based attacks against the same JBShield guard on llama-3.1-8b and compare them side-by-side with our EnumAttack.
Table 2. ASR(%) on llama-3.1-8b with JBShield (Zhang et al., 2025). While other attacks (Shen et al., 2024; Li et al., 2024a; Chang et al., 2024; Yong et al., 2024; Wei et al., 2023a) are effectively mitigated, JBShield is fully bypassed by our EnumAttack, which achieves even higher ASR by embedding the malicious query in the JSON schema payload. Caveats: IJP’s long prompts do not fit our schema (47.8% reflects template fit, not safety bypass); the SorryBench judge model does not understand Zulu/Base64, so their numbers are an LLM-as-judge artefact. Subsequent evaluations use gpt-4o as judge (see §§ 5.1).
Method IJP(Shen et al., 2024) DrAttack(Li et al., 2024a) Puzzler(Chang et al., 2024) Zulu(Yong et al., 2024) Base64(Wei et al., 2023a)
@@
As shown in Table 2, JBShield detects prompt-based attacks with high accuracy, reducing the ASR to nearly 0% regardless of the original attack strength, which shows its effectiveness against traditional data-plane threats. However, JBShield fails to detect EnumAttack, producing identical results with and without the guard. This occurs because JBShield is a lightweight defense that relies on static attack patterns and model-specific white-box prompt states. By keeping the prompt body fully benign, CDAs such as EnumAttack bypass this detection path, which makes static, data-plane-only defenses ineffective against control-plane attacks.
@@
To ensure our findings are not specific to a particular constrained-decoding implementation, we evaluate two grammar engines (xgrammar (Dong et al., 2025) and outlines (Willard and Louf, 2023)) across mainstream serving stacks (vllm (Kwon et al., 2023) and sglang (Zheng et al., 2024)). As shown in Figure 5, EnumAttack remains consistently successful across all combinations, indicating the vulnerability is inherent to the constrained-decoding mechanism rather than a specific implementation flaw. The minor variation under outlines stems from its less robust grammar handling, which occasionally produces empty or repeated-whitespace outputs counted as failed attacks.
Figure 5. EnumAttack ASR across grammar backends (xgrammar, outlines) and serving engines (vllm, sglang) on llama-3.1-8b. Apart from negligible variation, the attack is consistently successful across all combinations.
@@
Table 4. Major ASR results on mainstream deployed models (gpt-4o, gpt-4o-mini, gemini-2.0-flash). Results w/ and w/o grammar guardrail are listed outside/inside parentheses. We use llama-guard-3-8b as the grammar guardrail and set k=1k{=}1 for DictAttack.
[…]
Baseline 2.1%(2.1%) 44.0%(44.0%) 14.0%(14.0%) 42.5%(42.5%) 8.9%(8.9%)
gpt-4o-mini EnumAttack 3.4%(99.0%) 7.0%(100.0%) 8.0%(97.0%) 16.8%(96.4%) 5.1%(98.1%)
[…]
Model AdvBench HarmBench JailbreakBench SorryBench StrongREJECT
5.3. CDAs Resist Defenses Built to Audit the Control Plane Itself
When defenders move onto the control plane — auditing the schema, or jointly auditing prompt and schema with a Dual-Plane Guard — CDAs remain evasive. DictAttack still achieves 75.8% ASR against the strongest such defense. The cross-plane decoupling we engineer in DictAttack keeps the malicious intent legible only to the model, not to the auditor. We turn on the grammar guard and a Dual-Plane Guard that audits prompt and schema jointly, and evaluate along two dimensions:
  1. (i)
  2. (ii)
5.3.1. Why Data-Plane Attacks Get Suppressed Here, but DictAttack Does Not
The Dual-Plane Guard audits both the prompt and the schema. For PAIR (Chao et al., 2024b), TAP (Mehrotra et al., 2024), and AutoDAN-Turbo (Liu et al., 2025), the entire malicious payload sits on the prompt side, so a stronger LLM auditor on that side easily catches them. DictAttack instead splits the payload across prompt and schema — neither half is harmful in isolation — so the same auditor cannot see the intent before generation. We measure this on gpt-4o over AdvBench: prompt-based attacks face the strongest baseline guard; DictAttack faces our Dual-Plane Guard.
Table 3 confirms this surface gap. AutoDAN-Turbo’s ASR drops from 96.0% (undefended) to 26.0% once the prompt-side auditor sees its payload; PAIR/TAP go from 11.5%/24.4% to 1.6%/0.4% for the same reason. DictAttack, whose harmful intent has no prompt-side or schema-side “shape” for the auditor to catch, keeps 75.8% ASR under the same Dual-Plane Guard. A finer-grained breakdown across $k$ and moderation backends is in §§§ 5.3.3.
5.3.2. Output Quality under StrongREJECT
ASR captures whether an attack succeeds, not whether the resulting answer is actually useful to an adversary. To address this gap, we additionally report the StrongREJECT (Souly et al., 2024) composite score, which combines success with response usefulness:
$\mathrm{SR}=\mathrm{ASR}\times\tfrac{1}{2}\bigl(\mathrm{convincing}+\mathrm{specific}\bigr),$
where convincing and specific are LLM-judged sub-scores in $[0,1]$ measuring whether the response is fluent and operationally specific. We evaluate on gemini-2.5-pro over AdvBench, comparing DictAttack to three SOTA prompt-based jailbreaks under the same judge pipeline as §§§ 5.3.1.
Table 6. StrongREJECT (Souly et al., 2024) composite score on gemini-2.5-pro over AdvBench. Higher means a successful jailbreak is also more useful (more convincing and specific) to the attacker. DictAttack produces near-perfect-quality harmful answers, while typical prompt-based baselines either trigger safe-but-evasive completions (PAIR, TAP) or partially compromise quality.
Attack Method StrongREJECT score ($\uparrow$)
PAIR (Chao et al., 2024b) 0.10
TAP (Mehrotra et al., 2024) 0.23
AutoDAN-Turbo (Liu et al., 2025) 0.87
[…]
As shown in Table 6, DictAttack achieves a near-perfect composite of 0.98, indicating that its successful responses are both highly fluent and operationally specific. PAIR (0.10) and TAP (0.23) suffer because their adversarial prompts often coerce the model into evasive answers (e.g., generic non-actionable text) that pass a binary success check but score poorly on specific. AutoDAN-Turbo (0.87) is a strong baseline but, as shown in Table 3, remains substantially more vulnerable to data-plane guardrails than DictAttack. Across our evaluation we observe a consistent pattern: more capable models, once jailbroken via the control plane, generate more dangerous and higher-quality harmful content.
@@ 5.3.3. EnumAttack and DictAttack on grammar-guarded models
Table 4 and Table 5 report ASRs across five benchmarks. The baseline uses each benchmark’s original prompts, while EnumAttack and DictAttack wrap the same intents using our schema-based approach. For each entry, the value outside parentheses is the ASR with the grammar guard enabled and the value in parentheses is the ASR without the grammar guard. Baseline columns also reaffirm that direct prompting is largely mitigated by modern alignment (avg. 6.7–38.7%).
EnumAttack achieves an average ASR of 95.8% undefended, but simple grammar auditing reduces its success rate to 7.9% (Table 4). This confirms that its reliance on explicit literals makes it highly detectable, motivating the need for more evasive CDAs.
DictAttack remains highly effective even under grammar auditing. The same grammar-level defense is almost entirely ineffective against DictAttack. As shown in Table 4 and Table 5, DictAttack maintains high ASRs across both mainstream deployed models and flagship/high-capability models. For example, averaged across the five benchmarks, DictAttack achieves 94.3–99.5% ASR on gpt-5, gemini-2.5-pro, deepseek-r1, and gpt-oss-120b without grammar guard (inside parentheses); applying llama-guard-3-8b with k=1k{=}1 (1.7% detection rate) yields 92.7–97.8% ASR with guard (outside parentheses). This demonstrates that the vulnerability is not a flaw of older checkpoints but a fundamental issue that persists in the latest, most heavily aligned models. The slightly lower ASR for gpt-4o-mini (86.4%) is attributed to its reduced reasoning capability, which sometimes fails to correctly reconstruct the query from the dictionary rather than any safety-related refusal.
@@ 5.3.4. DictAttack with combined prompt- and grammar-guarding
Thus, we further evaluate a Dual-Plane Guard setting, adapted from the SOTA SelfDefend (Wang et al., 2025). We adapt SelfDefend’s audit prompt to take both the data-plane prompt (keys) and the control-plane grammar (dictionary), using gpt-4o as the auditing LLM.
@@
First, the results show that even unified auditing cannot stop DictAttack. llamaguard detects only 28.6% of attacks, while Moderation APIs detect almost none, likely due to limited model capacity. Even the SOTA SelfDefend (Wang et al., 2025), using its best strategy, reaches only an 18.5% detection rate on AdvBench. This is because modern structured schemas are too complex for LLMs to parse and audit reliably across planes, creating a reasoning gap that current auditing-based defences cannot bridge.
Second, increasing the synthesis factor $k$ markedly weakens the detection. Under Defense 1, gpt-4o attains $\approx 18\%$ detection rate at $k=0$; however, when $k$ is raised to 10, detection drops to 5.3%, as the malicious tokens are further diluted within the benign dictionary. This reveals that DictAttack exploits a fundamental blind spot: the lack of cross-plane coordination in LLM safety architectures.
Table 7. Detection rate on AdvBench (Chen et al., 2022) under Defense 1 (grammar-only) and Defense 2 (Grammar + Prompt Dual Guard, adapted from SelfDefend; gpt-4o auditor). This table reports detection rate only; since our guardrails are decoupled from the underlying LLM, the corresponding guarded ASR follows deterministically as $\mathrm{ASR}_{\mathrm{guard}}=(1-\mathrm{Detection})\times\mathrm{ASR}_{\mathrm{undef}}$. e.g., on gemini-2.5-pro (AdvBench, k=1k{=}1): $(1-1.7\%)\times 99.2\%\approx 97.5\%$, matching the value reported in Table 5.
@@ 5.3.5. Discussion: Space-Time Decoupling via Interleaved DictAttack
In this variant, the attacker splits the two DictAttack payloads across turns: the prompt payload (keys) is sent first and stored in the conversation context (KV cache), while the grammar payload (dictionary) arrives later in a separate request. Since practical dual-plane guards operate on a per-request basis, this temporal decoupling collapses dual-plane auditing into a single-plane check, resulting in only 5.3% detection for gpt-4o.
Mitigating this would require retrieving and auditing the entire conversation history at every turn, which is impractical. First, the cost would be prohibitive because the guard cannot know in advance when a malicious payload will appear, so it would need to repeatedly scan long histories. Second, as the context grows toward gpt-4o’s 128k-token limit, the guard would need context length and reasoning capacity comparable to or exceeding that of the protected model, making real-time deployment infeasible. Finally, modern JSON schemas are dense and opaque, so reconstructing a hidden query from a dictionary buried dozens of turns back becomes a “needle in a haystack” problem that current guard models are not built to solve.
In conclusion, DictAttack demonstrates that by decoupling intent across both space (planes) and time (requests), adversaries can render external auditing defenses ineffective and expensive, and ultimately impractical.